A Company placed into service an apartment building in March, 2018 which costs $200,000. Using MACRS show the depreciation for 2018, 2019 and 2020. Show the adjusted basis after the 2020 depreciation
Answer #1

Depreciation Rates

Year 3 Years
1 33.33 %
2 44.45 %
3 14.81 %
4 7.41 %
YEAR DEPRECIATION RATE DEPRECIATION
2018 33.33 % $200000 * 33.33/100 = $ 66,660
2019 44.45 % $200000 * 44.45/100 = $ 88,900
2020 14.81 % $ 200000 * 14.81/100 = $ 29.620
2021 7.41% $ 200000* 7.41/100 = $ 14,820